# Functional vs. Object-Oriented: Comparing How Programming Paradigms Affect the Architectural Characteristics of Systems

### v1

@@ Abstract.
After decades of dominance by object-oriented programming (OOP), functional programming (FP) is gaining increasing attention in the software industry (Finley, 2022). This study compares the impact of OOP and FP on the architectural characteristics of software systems. For that, it examines the design and implementation of a Digital Wallet system, developed in Kotlin (representing OOP) and Scala (representing FP). The comparison is made through both qualitative and quantitative analyses to explore how each paradigm influences the system’s architectural characteristics. The self-ethnographic qualitative analysis provides a side-by-side comparison of both implementations, revealing the perspective of those writing such code. The survey-based quantitative analysis gathers feedback from developers with diverse backgrounds, showing their impressions of those reading this code. Hopefully, these results may be useful for developers or organizations seeking to make more informed decisions about which paradigm is best suited for their next project.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Brazilian Symposium on Software Quality; Nov 04–07, 2025; São José dos Campos, SPcopyright: noneccs: Software and its engineering Functional languagesccs: Software and its engineering Object oriented languagesccs: Software and its engineering Software usabilityccs: Software and its engineering Software reliabilityccs: Software and its engineering Software fault tolerance
@@ 1. Introduction
After decades of dominance of Object Oriented Programming (OOP), the Functional Programming (FP) paradigm has gained significant attention in the software industry (Finley, 2022). This shift is a response to growing complexities in software systems, the demand for better scalability, and the need for more predictable, robust, and maintainable codebases.
According to the IEEE Spectrum magazine in 2022 (Scalfani, 2022), one of the clearest indications of this shift is the increasing incorporation of functional features into different mainstream programming languages. Evaluating how FP compares to OOP is essential to understand the future of software development.
This research aims to contribute with insights into how OOP and FP influence the architectural characteristics of software systems. Hopefully, this can help developers and organizations deciding on which paradigms is best suited for their projects.
2. Research Questions
@@
RQ1: How do the functional and object-oriented paradigms impact different architectural characteristics of a system?
According to Richards and Ford (Richards and Ford, 2020), an architectural characteristic describes a concern that follow three criteria:
@@
    It specifies a non-domain design consideration.
@@
    It influences structural aspects of the design.
@@
    It is critical or important to the application success.
Section˜2 aims to understand the differences that arise when the same system is developed under different paradigms. In particular, this research focused on six architectural characteristics, summarized in table˜1. They encompass concerns that highlight important differences between the paradigms, facilitating the evaluation of their strengths and limitations in various aspects of software development.
@@
During the capstone project this research is based on, section˜2 was divided in five research sub-questions based on the architectural characteristics from table˜1. Considering space constraints, this paper focuses on only three of them, whose answers brought the most interesting discussions detailed in section˜9.
RQ1.1: How do functional and object-oriented paradigms impact the extensibility of a system?
@@
Section˜2 aims to explore how each paradigm influences the ability to add new components to the system. OOP typically relies on object composition and class inheritance, while FP uses high-order functions and generic data structures.
RQ1.2: How do functional and object-oriented paradigms impact the reusability of a system?
@@
Section˜2 aims to explore how each paradigm influences the ability to reuse components in the system. OOP typically relies on design patterns that take advantage of polymorphism, while FP combines list-based immutable data types with closures and recursion.
@@
RQ1.3: How do the functional and object-oriented paradigms impact error handling / propagation of a system?
@@
Section˜2 aims to explore how each paradigm embodies distinct philosophies in error management. OOP typically uses exceptions, while FP applies monadic types.
@@ 3. Methodology
This section details the research methodology employed in this project, whose goal is to answer section˜2.
@@ 3.1. Proof of Concept
The object-oriented version of the Digital Wallet system was based on the Kotlin programming language, which supports classic object-oriented principles, such as abstraction, encapsulation, polymorphism, and inheritance. The developer can define classes in a straightforward manner, enabling the structuring of data and behavior through objects.
@@
The functional version of the Digital Wallet system was based on the Scala programming language, which supports classic functional programming constructs, such as higher-order functions, immutable data structures, pattern matching, and monads. The developer can focus on immutability and pure functional constructs, aligning with the core principles of functional programming.
Section˜4 describes the set of functional requirements that guided the development of both versions of the Digital Wallet system. Sections˜5 and 6 describe the main business logic that was implemented in both proof of concepts. The code developed during the research is available in the paper’s reproduction package.
@@ 3.2. Analysis
The second phase of the research consists of an analysis of the paradigms using a mixed-method approach. By triangulating benefits and drawbacks of each paradigm over the architectural characteristics, it is possible to provide an answer to section˜2.
The analysis will be based on the implementation of the Digital Wallet system. In particular, the focus is on pairs of functionally equivalent code snippets, i.e., code snippets that perform the same task, from both proofs of concept.
@@
It represents the perspective of writing code in the different paradigms. In particular, the self-ethnographic methodology means that it presents the assessments of the authors of the paper, who developed the proofs of concept. These results are further explored in section˜7.
@@
It represents the perspective of reading code in the different paradigms. In particular, the survey-based methodology means that it summarizes the assessments of different developers with varying levels of experience with the paradigms. These results are further explored in section˜8. The questionnaire and its results are available in the paper’s reproduction package.
After presenting these results, section˜9 combines these perspectives to answer the research subquestions.
@@ 4.1. Wallet Management
It represents a portfolio of investment options where customers can allocate their funds. Funds are invested by transferring funds from the Real Money Wallet to the Investment Wallet, and are liquidated the other way around. Investments and liquidations can be initiated at any time but will only settle on the next business day.
@@ 4.2. Transaction Lifecycle Management
[…]
Figure 1. Transaction Lifecycle.
@@
Figure˜1 summarizes the compatibility between transactions and different wallets. Every transaction must be validated before execution. Deposit and Withdrawal transactions are restricted to the Real Money Wallet. Transfer transactions are allowed only between Real Money Wallet and Emergency Funds Wallet. Hold transactions can be placed only on Real Money Wallet and Investment Wallet, whereas Transfer From Hold transactions are limited to be executed between them. Finally, all transactions – except for Deposit – require a balance check to ensure sufficient funds are available to complete the operation.
The Digital Wallet system must manage transaction failures as part of the transaction lifecycle. Transactions failing validation should be marked as permanently failed, while those that fail during execution should be eligible for retry. In general, validation should capture expected, unrecoverable errors, such as insufficient funds. Execution errors, though unlikely after successful validation, may still occur due to internal issues or third-party API instabilities. In such cases, the system should provide a mechanism to retry the transaction.
@@ 4.3. Investment Customization
[…]
Example Investment Policy
Figure 2. Example Investment Policy. The investments are divided as follows: $50\%$50 % to stocks, $30\%$30 % to real estate, $15\%$15 % to bonds, and $5\%$5 % to cryptocurrency.
The Digital Wallet system should allow customizing investments and liquidations. The participation should be measured as a percentage of the total invested under an Investment Wallet, and the customer should have real-time control over it. This requirement includes the capacity of setting the percentage to zero for undesired investments or even selecting a single investment.
@@ 5.1. Wallet
The Wallet entity models an account where customer funds are held and managed. It functions as the hub for all financial transactions, allowing money to be transferred (deposited or withdrawn) to other destinations. The Digital Wallet system provides three types of wallets, as specified in section˜4.1: Real Money Wallet entity, Investment Wallet entity, and Emergency Fund Wallet entity.
While Transaction entities dictate the flow of funds between Wallets entities, the latter are unaware of the former’s existence. Their role is to model and hold customer funds within the system. They are not responsible for determining the validity of any incoming or outgoing Transaction entity. This separation of concerns ensures that Wallet entities focus on balance management without handling transaction logic.
@@ 5.2. Subwallet
[…]
Wallet entities and corresponding Subwallet entities.
Figure 3. Wallet entities and corresponding Subwallet entities. The Investment Wallet illustrates multiple Subwallet entities associated to a Wallet entity.
The Real Money Wallet has only one Subwallet entity: the Real Money Subwallet. Similarly, the Emergency Funds Wallet entity has only one Subwallet entity: the Emergency Funds Subwallet. In practice, having a single Subwallet entity means that, for record-keeping purposes, funds in the Wallet entity are not divided in any way.
The relevance of the Subwallet entity becomes evident in the Investment Wallet, where funds are allocated across various investment options, as illustrated by figure˜3. Introducing a new investment option involves composing new Subwallet entities within the Investment Wallet.
@@ 5.3. Transaction
A Transaction can be classified as a Deposit, Withdrawal, Hold, Transfer, or Transfer From Hold. Each transaction type represents a standard operation that modifies the state of a Wallet as specified in section˜4.2.
@@
A Transaction must be validated to ensure they adhere to the internal rules of the Digital Wallet system. For example, a Deposit and Withdrawal can only be initiated on a Real Money Wallet, but not on an Investment Wallet. These validations follow the restrictions specified in section˜4.2.
@@
A Transaction is created with the Processing status, indicating they have not yet gone through validation or execution. The Failed status is a terminal state, representing a permanent failure with no possibility of retry. Similarly, the Completed status is a terminal state, representing that all validations and execution steps succeeded, the transaction has settled, and the involved Wallet entity was updated. The Transient Error status is a cyclic state, representing a recoverable error occurred during transaction processing (such as a failure in a third-party API call), which is eligible for retry.
@@ 5.4. Journal Entry
A Journal Entry entity documents the effects of a Transaction on a Wallet. It tracks changes in the Wallet balance and provides a detailed history of all financial activities. By listing all Journal Entry entities linked to a Wallet, it is possible to calculate its current state accurately.
@@
  • wallet_id
    It identifies the source or target Wallet involved in the transaction. It may be absent to represent an external bank account for Deposit and Withdrawal transaction types.
  • subwallet_id
    It identifies the Subwallet under the main Wallet involved in the transaction. Similarly, it may also be absent to represent an external bank account.
  • amount
    It accounts for the money being added (if positive) or deducted (if negative) from the Subwallet.
  • balance_type
    It can have three values: Internal, for funds being sent to or received from an external bank account; Available, for funds being added to or deducted from the available balance of a Subwallet; and Holding, for funds being added to or deducted from the reserved balance of a Subwallet.
The Ledger entity collects all Journal Entry entity created by the Digital Wallet system. Each Transaction with Completed status results in a pair of Journal Entry being recorded in the Ledger. Figure˜4 illustrates how these journal entries look like for each transaction type.
[…]
Example of pairs of Journal Entry created for each Transaction type.
Figure 4. Example of pairs of Journal Entry created for each Transaction type.
@@ 5.5. Investment Policy
An Investment Policy entity defines the allocation of an investment portfolio by associating each Subwallet with a specific percentage representing its share of the total portfolio, as specified in section˜4.3.
@@ 6. Digital Wallet System: Services
The Digital Wallet system relies on a set of services that collectively operate the functionalities described in section˜4.
Figure˜5 summarizes the existing services and illustrates the communication flow between them.
[…]
Services that compose the Digital Wallet system.
Figure 5. Services that compose the Digital Wallet system.
@@ 6.1. Ledger Service
The Ledger service is responsible for intermediating the interactions between other services and the Digital Wallet system Ledger entity. It provides two functionalities: (1) posting new Journal Entry, and (2) querying the Ledger for its balance. For the latter, it aggregates Journal Entry to compute the balance of a specific Wallet.
@@ 6.2. Transaction Service
The Transactions service provides functionalities to fulfill each step of a transaction lifecycle specified in section˜4.2. It creates, validates, and processes Transaction s, encapsulating the business logic for each Transaction type.
@@ 6.3. Wallet Service
  • Deposit Request:
    Deposits funds into the Real Money Wallet.
  • Withdraw Request:
    Withdraw funds from the Real Money Wallet.
  • Emergency Allocation Request:
    Withdraw funds from the Real Money Wallet.
  • Emergency Release Request:
    Transfers funds from the Emergency Funds Wallet back to the Real Money Wallet.
  • Investment Request:
    Allocates funds from the Real Money Wallet to the Investment Wallet, distributing the amount across investment options based on the InvestmentPolicy.
  • Liquidation Request:
    Withdraws funds from the Investment Wallet to the Real Money Wallet, proportionally liquidating investments according to the InvestmentPolicy.
6.4. Investment Service

### v3

@@ ABSTRACT.
This study compares the impact of adopting object-oriented programming (OOP) or functional programming (FP) on the architectural characteristics of software systems. For that, it examines the design and implementation of a Digital Wallet system developed in Kotlin (for OOP) and Scala (for FP). The comparison is made through a mixed-method approach. The self-ethnographic qualitative analysis provides a side-by-side comparison of both implementations, revealing the perspective of those writing such code. The survey-based quantitative analysis gathers feedback from developers with diverse backgrounds, showing their impressions of those reading this code. Hopefully, these results may be useful for developers seeking to decide which paradigm is best suited for their next project.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Brazilian Symposium on Software Quality; Nov 04–07, 2025; São José dos Campos, SPcopyright: none
@@ 1. Introduction
After decades of dominance of Object-Oriented Programming (OOP), the Functional Programming (FP) paradigm has gained significant attention in the software industry (Finley, 2022). This shift is a response to growing complexities in software systems, the demand for better scalability, and the need for more predictable, robust, and maintainable codebases.
According to the IEEE Spectrum magazine in 2022 (Scalfani, 2022), one of the clearest indications of this shift is the increasing incorporation of functional features into different mainstream programming languages. Evaluating how FP compares to OOP is essential to understanding the future of software development.
@@
RQ: How do the functional and object-oriented paradigms impact different architectural characteristics of a system?
According to Richards and Ford (Richards and Ford, 2020), an architectural characteristic describes a concern that follows three criteria:
@@
    it specifies a non-domain design consideration,
@@
    it influences structural aspects of the design, and
@@
    it is critical or important to the application’s success.
This research focuses on six architectural characteristics, summarized in table 1. They encompass concerns that highlight important differences between the paradigms, making it easier to compare their strengths and limitations.
@@
Due to space constraints, this paper will address only four of the six architectural characteristics from table 1, whose analyses brought the most interesting results. For a full discussion, please refer to the bachelor thesis that originated this paper (Dias et al., 2024).
sRQ1: How do the functional and object-oriented paradigms impact the extensibility of a system?
@@
Section 1 aims to explore how each paradigm influences the ability to add new components to the system. OOP typically relies on object composition and class inheritance, while FP uses high-order functions and generic data structures.
sRQ2: How do the functional and object-oriented paradigms impact the reusability of a system?
@@
Section 1 aims to explore how each paradigm influences the ability to reuse components in the system. OOP typically relies on design patterns that take advantage of polymorphism, while FP combines list-based immutable data types with closures and recursion.
@@
sRQ3: How do the functional and object-oriented paradigms impact error handling and propagation of a system?
@@
Section 1 aims to explore how each paradigm embodies distinct philosophies in error management. OOP typically uses exceptions, while FP applies monadic types.
2. Programming Paradigms
Understanding the principles of OOP and FP is essential for a meaningful comparison between them. Their distinct philosophies, strengths, and limitations significantly influence how software systems approach architectural characteristics.
2.1. Object-Oriented Programming
According to Martin (Martin, 2017), any OOP language must support these three characteristics: encapsulation, inheritance, and polymorphism.
Encapsulation
It ensures that a clear boundary is established around a cohesive set of data and behaviors (Martin, 2017). Within this boundary, data is hidden and protected from direct external access, while only specific functions are exposed to interact with the data. This approach is commonly seen in object-oriented programming through private data members and public methods in classes.
Inheritance
It is the process of reusing and extending a group of variables and functions defined in one scope within another (Martin, 2017). In OOP, this often means that a new class (called a subclass) can inherit properties and behaviors from an existing class (called a superclass). This allows the new class to reuse, override, or expand existing functionality without redefining it entirely.
Polymorphism
It is the ability of a method to behave differently based on the context in which it is used (Martin, 2017). Polymorphism unblocks dependency inversion by allowing high-level modules to depend on abstractions rather than concrete implementations. It decouples implementation details from the modules that depend on them through interfaces, allowing dependencies to be reversed.
2.2. Functional Programming
According to Pilquist et al. (Pilquist et al., 2023), functional programming is based on the premise of writing programs using only pure functions, which leads to other patterns like high-order functions and monadic types.
Pure Functions
They are functions without side effects (Martin, 2017). A function has side effects if it does something other than returning a result, such as modifying a variable, throwing an exception, or printing to the console.
High-Order Functions
In FP, functions are values (Pilquist et al., 2023). They can be assigned to variables, stored in data structures, and passed as arguments to functions. A function that accepts other functions as arguments is called a higher-order function (HOF).
Monadic Types
Most FP languages support pattern matching (Pilquist et al., 2023) as a way to execute functions depending on the value passed to them. Constructs such as Option and Either have the properties of Monads (Pilquist et al., 2023), types that leverage pattern matching to provide clean composition and error handling.
3. Related Literature
Few works in the literature focus on the same research questions proposed in section 1. A search with the keywords “Functional versus Object-Oriented Programming” on on Google Scholar results in papers that propose to mix the paradigms in the context of a single programming language or application.
In the late 1990s, Harrison et al. (Harrison et al., 1996; Samaraweera and Harrison, 1998) compared the code of $12$ programs produced in SML (FP) and C++ (OOP), measuring their difference on development and code metrics. In 2016, Alic et al. (Alic et al., 2016) compared algorithms in Java and C# (OOP) versus Haskell and F# (FP), focusing on lines of code and efficiency. In 2025, Khan et al. (Khan et al., 2025) compared the impact of choosing OOP, FP or Procedural Programming for implementing microservices architectures (Newman, 2021), analyzing their impact on application design.
The current gap in the literature is an opportunity to explore the differences between paradigms once again, using modern programming languages and focusing on both code and design.
@@ 4. Methodology
This section details the research methodology employed in this project, whose goal is to answer section 1.
@@ 4.1. Proof of Concept
The object-oriented version of the Digital Wallet system was based on the Kotlin programming language, which supports the OOP principles from section 2.1. The developer can define classes in a straightforward manner, enabling the structuring of data and behavior through objects.
@@
The functional version of the Digital Wallet system was based on the Scala programming language, which supports classic FP principles from section 2.2. The developer can focus on immutability and pure functional constructs, aligning with the core principles of functional programming.
Section 5 describes the set of functional requirements that guided the development of both versions of the Digital Wallet system. Sections 6 and 7 describe the main business logic that was implemented in both proofs of concept. The code developed during the research is available in the paper’s reproduction package.
@@ 4.2. Analysis
The second phase of the research consists of an analysis of the paradigms using a mixed-method approach. By triangulating benefits and drawbacks of each paradigm over the architectural characteristics, it is possible to provide an answer to section 1.
The analysis will be based on the implementation of the Digital Wallet system. In particular, the focus is on code snippets that perform the same task from both proofs of concept.
@@
It represents the perspective of writing code in the different paradigms. In particular, the self-ethnographic methodology means that it presents the assessments of the authors of the paper, who developed the proofs of concept. These results are further explored in section 8.
@@
It represents the perspective of reading code in the different paradigms. In particular, the survey-based methodology means that it summarizes the assessments of different developers with varying levels of experience with the paradigms. These results are further explored in section 9.
After presenting these results, section 10 combines these perspectives to answer the research subquestions.
@@ 5.1. Wallet Management
It represents a portfolio of investment options where customers can allocate their funds. Funds are invested by transferring funds from the Real Money Wallet to the Investment Wallet and are liquidated the other way around. Investments and liquidations can be initiated at any time but will only settle on the next business day.
@@ 5.2. Transaction Lifecycle Management
Figure 1. Transaction Lifecycle. :
@@
Figure 1 summarizes the compatibility between transactions and different wallets. Every transaction must be validated before execution. Deposit and Withdrawal transactions are restricted to the Real Money Wallet. Transfer transactions are allowed only between Real Money Wallet and Emergency Funds Wallet. Hold transactions can be placed only on Real Money Wallet and Investment Wallet, whereas Transfer From Hold transactions are limited to being executed between them. Finally, all transactions – except for Deposit – require a balance check to ensure sufficient funds are available to complete the operation.
The Digital Wallet system must manage transaction failures as part of the transaction lifecycle. Transactions failing validation should be marked as permanently failed, while those that fail during execution should be eligible for retry. In general, validation should capture expected, unrecoverable errors, such as insufficient funds. Execution errors, though unlikely after successful validation, may still occur due to internal issues or third-party API instabilities. In such cases, the system should allow to retry the transaction.
@@ 5.3. Investment Customization
The Digital Wallet system should allow customizing investments and liquidations. The participation should be a percentage of the total invested under an Investment Wallet, and the customer should have real-time control over it. This requirement includes the capacity of setting the percentage to zero for undesired investments or even selecting a single investment.
@@ 6.1. Wallet
The Wallet entity models an account where customer funds are held and managed. It functions as the hub for all financial transactions, allowing money to be deposited or withdrawn to other destinations. The Digital Wallet system provides three types of wallets, as specified in section 5.1Real Money Wallet, Investment Wallet, and Emergency Fund Wallet.
While Transaction entities dictate the flow of funds between Wallets entities, the latter are unaware of the former’s existence. Their role is to model and hold customer funds within the system. They are not responsible for determining the validity of any incoming or outgoing Transaction. This separation of concerns ensures that Wallet entities focus on balance management without handling Transaction logic.
@@ 6.2. Subwallet
The Real Money Wallet has only one Subwallet entity: the Real Money Subwallet. Similarly, the Emergency Funds Wallet entity has only one Subwallet entity: the Emergency Funds Subwallet. In practice, having a single Subwallet means that, for record-keeping, funds in the Wallet are not divided in any way.
Figure 2. Wallets and corresponding Subwallets. :
The relevance of the Subwallet entity becomes evident in the Investment Wallet, where funds are allocated across various investment options, as illustrated by figure 2. Introducing a new investment option involves composing new Subwallet entities within the Investment Wallet.
@@ 6.3. Transaction
A Transaction can be classified as a Deposit, Withdrawal, Hold, Transfer, or Transfer From Hold. Each transaction type represents a standard operation that modifies the state of a Wallet as specified in section 5.2.
@@
A Transaction must be validated to ensure they adhere to the internal rules of the Digital Wallet system. For example, a Deposit and Withdrawal can only be initiated on a Real Money Wallet, but not on an Investment Wallet. These validations follow the restrictions specified in section 5.2.
@@
A Transaction is created with the Processing status, indicating they have not yet gone through validation or execution. The Failed status is a terminal state, representing a permanent failure with no possibility of retry. Similarly, the Completed status is a terminal state, representing that all validations and execution steps succeeded, the transaction has settled, and the involved Wallet entity was updated. The Transient Error status is a cyclic state, representing a recoverable error that occurred during transaction processing (such as a failure in a third-party API call), which is eligible for retry.
@@ 6.4. Journal Entry
A Journal Entry entity documents the effects of a Transaction on a Wallet. It tracks changes in the Wallet balance. By listing all Journal Entry entities linked to a Wallet, it is possible to calculate its current state accurately.
@@
  • wallet_id: It identifies the source or target Wallet. It may be absent to represent an external bank account.
  • subwallet_id: It identifies the Subwallet under a Wallet. It may also be absent to represent an external bank account.
  • amount: It accounts for the money being added (if positive) or deducted (if negative) from the Subwallet.
  • balance_type: It can have three values: Internal, for funds sent to or received from an external bank account; Available, for funds added to or deducted from the available balance of a Subwallet; and Holding, for funds added to or deducted from the reserved balance of a Subwallet.
The Ledger entity collects all Journal Entry entity created by the Digital Wallet system. Each Transaction with Completed status results in a pair of Journal Entry being recorded in the Ledger. Figure 3 illustrates how these journal entries look like for each transaction type.
Figure 3. Pairs of Journal Entry for each Transaction type. :
@@ 6.5. Investment Policy
An Investment Policy entity defines the allocation of an investment portfolio by associating each Subwallet with a specific percentage representing its share of the total portfolio, as specified in section 5.3.
@@ 7. Digital Wallet System: Services
The Digital Wallet system relies on a set of services that collectively operate the functionalities described in section 5.
Figure 4 summarizes the existing services and illustrates the communication flow between them.
Figure 4. Services that compose the Digital Wallet system. :
@@ 7.1. Ledger Service
The Ledger service is responsible for orchestrating between other services and the Digital Wallet system Ledger entity. It provides two functionalities: (1) posting new Journal Entry, and (2) querying the Ledger for its balance. For the latter, it aggregates Journal Entry to compute the balance of a specific Wallet.
7.2. Investment Service
@@ 7.3. Transaction Service
The Transactions service provides functionalities to fulfill each step of a transaction lifecycle specified in section 5.2. It creates, validates, and processes Transactions, encapsulating the business logic for each Transaction type.
@@ 7.4. Wallet Service
  • Deposit: Adds funds into the Real Money Wallet.
  • Withdraw: Removes funds from the Real Money Wallet.
  • Emergency Allocation: Adds funds into the Emergency Funds Wallet.
  • Emergency Release: Transfers funds from the Emergency Funds Wallet to the Real Money Wallet.
  • Investment: Transfers funds from the Real Money Wallet to the Investment Wallet, distributing the amount across investment options based on the InvestmentPolicy.
  • Liquidation: Transfers funds from the Investment Wallet to the Real Money Wallet, liquidating investments proportionally according to the InvestmentPolicy.